need help urgent!
1. It takes jun thrice the number of days it takes Peter to mow a lawn. When the two boys work together, the job is completed in three days. How long does it take each boy to do the job alone?

Answer by Alan3354(69443)

1. It takes jun thrice the number of days it takes Peter to mow a lawn. When the two boys work together, the job is completed in three days. How long does it take each boy to do the job alone?
------------------------
It takes Jun 3x # of days to do the job, and Peter x days.
So, Peter does 1/x of the job per day, and Jun does 1/3x of the job.
Working together, they do the sum, 1/x + 1/3x per day.
1/x + 1/3x = 4/3x per day.
4/3x jobs per day --> 3x/4 days to do the job together.
3x/4 = 3
3x = 12
x = 4
It takes Peter 4 days.
It takes Jun 12 days.